5) Consider an electron located on the x-axis such that the uncertainty in its position is 1.00 Angstrom. What is the minimum uncertainty in its momentum along the x axis? Give the answer in units of (kg m)/sec 6) If the work function of a metal is 2.00 x 10-19 J, what is the frequency of the light required to just eject an electron from the metal, that is, such that the ejected electron has zero kinetic energy? Give the answers in sec-1 .
Answer #1

Solution) We have (change in position)×(change in momentum)=h/(4×pi)

Let change in position be X

Change in momentum P

But X.P=h (approximately)

P=h/X

P=6.625×10^(-34)/(1×10^(-10))

P=6.625×10^(-24) kgm/s

For workfunction (phi)=2×10^(-19)J

Frequency f=?

We have phi =hf

f=phi/h

f=2×10^(-19)/(6.625×10^(-34)

f=0.30×10^(15) sec^(-1)